There are four people playing a game with a standard die. In the game, the person who rolls the lowest number wins. Ties are broken by continuing to roll until the tie is broken. One of the four people rolls a three. What is the probability that they will win?
Answer by stanbon(75887)

There are four people playing a game with a standard die. In the game, the person who rolls the lowest number wins. Ties are broken by continuing to roll until the tie is broken. One of the four people rolls a three. What is the probability that they will win?
----
The only way they can lose is for someone to roll a 1 or a 2.
---
The way they will win is for others to roll 4,5, or 6.
---
P(they will win) = 3/6 = 1/2
===================================
Cheers,
Stan H.